OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 2x^2 + 9y^2 = 18

OpenStudy (anonymous):

rewrite it in the form \[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] then you can get the rest

OpenStudy (anonymous):

Ohhh ok great

OpenStudy (anonymous):

Thanks so much